# Four special Poncelet triangle families about the incircle

### v1

@@ 1. Introduction
In the spirit of the research done in [garcia2024-incircle], we study four special families of ellipse-inscribed Poncelet triangles about the incircle. Referring to Figure 1, these are:
@@
  • iso-$X_{2}$: caustic is centered on a first special point such that the barycenter $X_{2}$ is stationary;
  • focal-$X_{4}$: caustic is centered on a second special point such that the orthocenter $X_{4}$ is stationary on a focus of $\mathcal{E}$;
  • iso-$X_{7}$: caustic is centered on a third special point such that the Gergonne point $X_{7}$ is stationary.
In the calculations below, $R$, $r$, $l_{i},i=1,2,3$ refer to a triangle’s circumradius, inradius, and sidelengths, respectively.
@@ 2. Focal-$X_{1}$
Referring to Figure 1 (top left):
@@
$r=\frac{b\sqrt{a^{4}-c^{2}x_{c}^{2}}-a\sqrt{b^{4}+c^{2}y_{c}^{2}}}{c^{2}}\cdot$
@@
For the focal-$X_{1}$ family, the invariant sum of half-angle sines is given by:
@@ 5. Iso-$X_{7}$
$|X_{1}-X_{7}|^{2}=r^{2}\left(1-\frac{3s^{2}}{(r+4R)^{2}}\right)\raisebox{1.29167pt}{.}$

### v2

@@ 1. Introduction
As special cases to [garcia2024-incircle], we introduce four special families of Poncelet triangles inscribed in an ellipse $\mathcal{E}$ and circumscribing their fixed incircle, let $C$ be its center, i.e., the incenter $X_{1}$ (triangle center notation $X_{k}$ are after Kimberling [etc]). Referring to Figure 1, the four Poncelet families are:
@@
  • iso-$X_{2}$: caustic is centered on a special point on $\mathcal{E}$’s minor axis such that the barycenter $X_{2}$ is stationary;
  • focal-$X_{4}$: caustic is centered on a special point $\mathcal{E}$’s major axis such that the orthocenter $X_{4}$ is stationary on a focus of $\mathcal{E}$;
  • iso-$X_{7}$: caustic is centered on a another special point on $\mathcal{E}$’s major axis such that the Gergonne point $X_{7}$ is stationary.
In the sections below, $R$, $r$, $l_{i},i=1,2,3$ refer to a triangle’s circumradius, inradius, and sidelengths, respectively, see [mw] for definitions.
@@ 2. Focal-$X_{1}$
Let $\mathcal{E}$ be given by $(x/a)^{2}+(y/b)^{2}=0$. Referring to Figure 1 (top left):
@@
$r=\frac{b\sqrt{a^{4}-c^{2}x_{c}^{2}}-a\sqrt{b^{4}+c^{2}y_{c}^{2}}}{c^{2}}\cdot\qed$
@@
For the focal-$X_{1}$ family, the sum of half-angle sines is invariant and given by:
@@ 5. Iso-$X_{7}$
$|X_{1}-X_{7}|^{2}=r^{2}\left[1-\frac{3s^{2}}{(r+4R)^{2}}\right]\raisebox{1.29167pt}{.}$